OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = -2x4 + 4x3 + 3x2 + 18

ganeshie8 (ganeshie8):

find all the factors of contant term, 18

ganeshie8 (ganeshie8):

f(x) = -2x4 + 4x3 + 3x2 + 18 ^

ganeshie8 (ganeshie8):

factors of \(18 = \pm 1, \pm 2, \pm 3, \pm 6, \pm 9, \pm 18 \)

ganeshie8 (ganeshie8):

next, find all factors of leading coefficient f(x) = -2x4 + 4x3 + 3x2 + 18 ^

ganeshie8 (ganeshie8):

factors of \(2 = \pm ,1 \pm 2\)

ganeshie8 (ganeshie8):

possible rational zeroes are the ratios of both

OpenStudy (anonymous):

so all the factors and their ratios? or just the ratios?

ganeshie8 (ganeshie8):

just the ratios

ganeshie8 (ganeshie8):

possible rational roots = \(\large \pm \frac{1}{1} , \pm \frac{2}{1} ,\pm \frac{3}{1} ,\pm \frac{6}{1} ,\pm \frac{9}{1} ,\pm \frac{18}{1} \) and also, \(\large \pm \frac{1}{2} , \pm \frac{2}{2} ,\pm \frac{3}{2} ,\pm \frac{6}{2} ,\pm \frac{9}{2} ,\pm \frac{18}{2} \)

ganeshie8 (ganeshie8):

simplify the fractions, and remove any duplicates

OpenStudy (anonymous):

oh ok. thank you!

OpenStudy (anonymous):

that was really helpful!

ganeshie8 (ganeshie8):

np :) glad it helped !